OpenStudy (anonymous):

A car rental company charges 35 dollars a day plus 25 cents a mile to rent a car. what would be the cost to rent a car for one day and drive it 200 miles.

OpenStudy (anonymous):

85 dollars. Do you understand how I got that?

OpenStudy (anonymous):

No not at all

OpenStudy (anonymous):

Okay, well. Since it costs 35 dollars a day and 25 cents a mile and you only rent the car for one day and drive it 200 miles you can figure it out multiplying 200 by .25 to convert that into dollars. Then you add that amount to the 35 dollars. 35+50=85. Do you understand now?

OpenStudy (anonymous):

Ind of Lawls yeah

OpenStudy (anonymous):

So you could answer the problem your self if different numbers were presented?

OpenStudy (anonymous):

I dont know im sure ill have some more questions about it later tho

OpenStudy (anonymous):

Well, I'll try to explain it so that you may be able to do it yourself.

OpenStudy (anonymous):

With problems like these, you are presented with a down payment, the 35 dollars. And then you are given an amount that you are to pay based off of a set distance, the .25 cents per mile for 200 miles. So, the equation would look like this: 35 + 200(.25) = x.

OpenStudy (anonymous):

So, any other problem of the like, you would just have to add. If they rearranged the data, so that you drove for 3 days and 100 miles, the equation would be: 35(3)+100(.25)=x

OpenStudy (anonymous):

So, when you are doing problems such as these, the base equation would be: 35 times the amount of days + .25 cents times the amount of miles = the total.

OpenStudy (anonymous):

Now do you understand?